OpenStudy (xmoses1):

Evaluate |-6| + 2|-4| - 3|6|......

OpenStudy (anonymous):

|-6| + 2|-4| - 3|6| = 6+2(4)-3(6)= 6+8-18=14-18=-4 @Xmoses1

OpenStudy (anonymous):

-4

OpenStudy (xmoses1):

Oh there is multiplication in there